OpenStudy (anonymous):

find the real solutions of the equation. 4(x+1)^2+14(x+1)+6=0

OpenStudy (cruffo):

let t = x+1 then 4(x+1)^2+14(x+1)+6=0 is just 4t^2 + 14t + 6 = 0 note that 4*6 = 24 = 2*12 and 2+12 = 14, thus 4t^2 + 14t + 6 = 0 can be factored (by grouping) 4t^2 + 2t + 12t + 6 = 0 2t(2t + 1) + 6(2t + 1) = 0 (2t + 1)(2t + 6) = 0 so 2t + 1 = 0 which means that t = -1/2 and 2t+6 = 0 which means that t = -3 Now since t = x+1 we have that x + 1 = -1/2 which means that x = -3/2 and x + 1 = -3 which means that x = -4 To check: 4(-3/2+1)^2+14(-3/2+1)+6 4(-1/2)^2 + 14(-1/2) + 6 4(1/4) -7 + 6 1-7+6 0 check :) 4(-4+1)^2+14(-4+1)+6 4(-3)^2 + 14(-3) + 6 4(9) -42 + 6 36 - 42 + 6 0 check :) So the two solutions are x = -3/2 and x = -4

OpenStudy (anonymous):

Substitute (x+1) by a for example. The equation becomes \[4a ^{2}+14a+6=0\]